OpenStudy (anonymous):

what is the molar concentration of [I-] from the mixing of .02M Na2S2O3, 0.3M KI, and 0.1H2O2

OpenStudy (aaronq):

you need to be given volumes for the solutions.

OpenStudy (anonymous):

I have the volumes: Na2S2O3- .001L KI- .001L H202- .003L

OpenStudy (aaronq):

find the moles of \(I^-\), then find the new molarity taking into account the total volume (of the other solutions).

OpenStudy (anonymous):

How can I find the moles of I- with this information?

OpenStudy (aaronq):

use the molarity formula \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

Would I first use this formula to calculate the moles of the ions I3- and S2O3^2- and then add those together to get the resulting moles for I-. Sorry if this is confusing you.

OpenStudy (aaronq):

there is no need to use \(S_2O_3^{2-}\), the question is asking about \(I^-\), no? you find the moles of \(I^-\), then use the molarity equation with the total volume.

OpenStudy (anonymous):

Wait... I think that I get it. Once I find the moles for I3- I can use it to find the \[I ^{-}\] because of the equation \[3 I ^{-}(aq) + H2O2(aq) + 2 H3O ^{+}(aq)--- I3^{-} (aq) + 4 H2O(l)\]

OpenStudy (anonymous):

and then find the molarity

OpenStudy (aaronq):

i'm afraid you're overcomplicating the question. \(M_{I^-}=\dfrac{n_{I^-}}{L_{solution}}=0.3 M=\dfrac{n_{I^-}}{0.001L}\rightarrow n_{I^-}=0.0003~moles\) new molarity \(M_{I^-}=\dfrac{n_{I^-}}{L_{solution}}\rightarrow M_{I^-}\dfrac{0.0003~moles}{0.001~L+~0.001~L+0.003~L}\)

OpenStudy (aaronq):

there was supposed to be an equal sign in the last part \(M_{I^-}=...\)

OpenStudy (anonymous):

I know, I'm trying to fill out this table for a rate law and activation energy lab. I understood everything until I got to this part, but thank you for the help